Given below is a differential operational amplifier in block form and its internal transistor circuit with VCC=20V, VEE= 12 V , RC=5kΩ,RE=1kΩ and R1=2kΩ. Assume that the transistors are identical and have a βac=βdc=100. An ac signal Vi(t)=5sin(10t) is applied to the input terminals where Terminal 1 is Positive (non-invertinq) & Terminal 2 is negative (inverting) as stated in the cases below. Draw the output signals in each case.

: a) V1=Vi(t) , V2=Grounded. Vo(t)=Vo2-Vo1 )

b) V1=Grounded , V2=Vi(t). Vo(t)=Vo2-Vo1

c) V1=Vi(t) , V2=Grounded. Vo1(t)=Vo1 & Vo(t)=Vo2. (6 Points) d) V1=Grounded , V2=Vi(t) Vo1(t)=Vo1 & Vo(t)=Vo2.

1 INPUTS DA OUTPUTS 2 2 Fig. 2 (ii) +Vcc Rc RC output input 1 (+) Q1 Q2 input 2 O RE A RE R1 -VEE
Answer #1

-Kec-20 w ww 5k esk eo VA (+ + Vee le LvBe fik Blk sak 19v VEET 5k 35k Dic De iezt es {re ube + yie, beare ERE=lk Elk QR = 4kiel (re+ 5*) = 12 recCRETOR, recc 5k - ieg. (re +ək)=209 - Tel. (5k) = 0 fez (5x) - 130 a) 2, = 0, (+) = 5 Sin (not). V, -0 ivoi - ich Rc : -5kici icioiel Voje -5k (IMA) - 5 Sin(lot) v Vog - 0 = D Volt) = Vog - VOI = 0-65 Sin Ciot)) hsin 107 (b) :V,=Voi - - OV Volt)= Voz - VOL -5 Sin(100)-0 --5 Sindiot) v (C) V=V7Ct). Ve = 0 iez - 12 sk =0 Vog = -legre o Volt) -VO! Voi = -(d) V;=0 Vg - V;[t) Vi le lez Vq 5k 5k = 0 FIMA Vor-le, Rc Vor Q -leq Rc. - - (IMA)5K --5SIN (107M